### v2

# Training-Free Label Space Alignment for Universal Domain Adaptation

Abstract
Universal domain adaptation (UniDA) transfers knowledge from a labeled source domain to an unlabeled target domain, where label spaces often differ and the target domain often contains private classes. Previous UniDA methods primarily focused on visual space alignment, however, severe domain shifts often lead to visual ambiguities, undermining their ability to predict the label space shift. Our preliminary experiment shows that focusing exclusively on label space alignment can be a more effective solution to these limitations. While generative vision-language models (VLMs) offer strong zero-shot capabilities that could help uncover target-private classes, naively employing them introduces challenges such as noisy and semantically ambiguous labels. To address these challenges, we propose a novel approach that carefully utilizes generative VLMs to improve label space alignment. We introduce an adaptive thresholding strategy that leverages meaningful relationships between source and discovered target labels. This allows a training-free mechanism to filter out noisy and ambiguous labels, effectively refining the label space by identifying shared and discovering target-private label subsets. We then build a universal classifier for robust self-training, which effectively utilizes reliable pseudo-labels derived from the identified label spaces. The results reveal that the proposed method considerably outperforms existing UniDA techniques across key DomainBed benchmarks, delivering an average improvement of +7.9% in H-score and +6.1% in H3-score. Furthermore, incorporating self-training further enhances performance and achieves an additional (+1.6%) increment in both H- and H3-scores.
keywords:
Universal Domain Adaptation , Vision-Language Models , Foundation Models , Robustness , Self-Training
journal: Engineering Applications of Artificial Intelligence
[…]
[inst1]organization=Department of Artificial Intelligence, Korea University,addressline=145 Anam-ro, Seongbuk-gu, city=Seoul, postcode=02841, country=South Korea
[…]
[inst2]organization=OMRON SINIC X,addressline=5-24-5 Hongo, Bunkyo-ku, city=Tokyo, postcode=113-0033, country=Japan
1 Introduction
Domain shift, the discrepancy between training and real-world data, remains a pivotal challenge for learning generalizable vision-language representations. Unsupervised domain adaptation (UDA) (Long et al., 2015; Ganin et al., 2016; Saito et al., 2018) aims to transfer knowledge from a labeled source domain to an unlabeled target domain, based on the stringent assumption that the label space remains unchanged across domains. However, in practical settings, the target label space often comprises only a subset of source classes that overlap with both domains or include target-exclusive classes unique to the target domain. Various studies have explored these complexities. Open-set domain adaptation (ODA) and partial-set domain adaptation (PDA) assume that the label space of each domain is a subset of another’s (Panareda Busto and Gall, 2017; Cao et al., 2018). Nonetheless, these approaches rely on prior knowledge of label space shifts, which limits their real-world applicability. To address this constraint, Universal Domain Adaptation (UniDA) was introduced (You et al., 2019a; Saito et al., 2020) for broader applicability.
Figure 1: Comparison between CLIP Prompt, a zero-shot classifier enhanced with corresponding label text prompts, and previous state-of-the-art (SOTA) methods. Simply adding target-private classes or removing source-private classes leads to substantial gains in H-score, achieved entirely without additional training.
Figure 2: Examples of semantic ambiguity in generative VLM outputs: (a) noisy response, (b) hyponym prediction, and (c) synonym prediction.
In UniDA, a primary challenge is distinguishing target-private samples from source-private ones. Many studies have proposed innovative detection criteria utilizing entropy (Saito et al., 2020; Qu et al., 2023a), binary classifiers (Saito and Saenko, 2021; Lu et al., 2024), optimal transport (Chang et al., 2022), energy-based methods (Chen et al., 2022a), class prototypes (Li et al., 2021a; Lai et al., 2023), and calibration techniques (Deng and Jia, 2023). These methods still suffer from severe domain shifts, which involve not only style variations but also content changes among shared classes. The reason is that the methods rely on the geometric structure of visual embeddings; however, such features are often drastically shifted across domains, leading to inconsistent decisions.
This study addresses these limitations by focusing exclusively on label space alignment. To achieve this, we leverage the robust zero-shot capabilities of Vision-Language Models (VLMs) such as CLIP (Radford et al., 2021) and ALIGN (Li et al., 2021b). For instance, CLIP effectively recognizes diverse classes across domains by prompting label names through its fixed text encoder. To evaluate this ability, we conducted a pilot study assessing the performance of a zero-shot classifier across various category spaces within the UniDA framework. As shown in Fig. 2, zero-shot classifiers using customized label spaces exhibit substantial improvements in UniDA performance. Interestingly, merely augmenting the zero-shot classifier with oracle category spaces already surpasses existing UniDA approaches, most of which attempt to achieve alignment solely in the visual space. These findings suggest that addressing the UniDA problem is more efficiently achieved by aligning label spaces rather than visual spaces across domains.
Inspired by Fig. 2, we aim to identify target-private classes and obtain an empirical version of these classes, termed predicted target-private labels, by employing generative VLMs (Li et al., 2022, 2023). In particular, predicted target-private labels are derived from labels discovered in target images using generative VLMs. Nevertheless, distinguishing whether discovered labels belong to part of the source labels or not becomes challenging due to noise (Fig. 2a) and semantic ambiguities, such as hyponyms (Fig. 2b) and synonyms (Fig. 2c). For example, in Fig. 2b, the generative VLM outputs “Panasonic,” a semantic hyponym of “Monitor” in the source labels. Without proper alignment, “Panasonic” is incorrectly predicted as a target-private label, leading to degraded performance.
To address this, we introduce Training-free Label Space Alignment for UniDA (TLSA). Specifically, we first identify the raw labels of the target samples using a generative VLM. Subsequently, we implement three steps to eliminate incorrect or semantically ambiguous discovered labels from the source labels and accurately predict the true target private labels: (i) Synonym label alignment (Sec. 3.3): We identify synonyms in raw discovered labels using WordNet and align them with the source labels to remove synonyms; (ii) Semantic label alignment (Sec. 3.4): To resolve the ambiguous similarity between source and discovered labels, we identify a set of predictions that are considered correct for an image within the CLIP joint embedding space through an adaptive thresholding process. If the resulting set excludes source labels and includes only discovered labels, the filtered discovered labels are retained as candidate labels; (iii) Frequency-based noise-candidate filtering (Sec. 3.5): To eliminate residual noise, we monitor the frequency of target private candidate labels that pass previous filters across the entire target domain, removing noisy candidate labels based on their occurrence patterns.
Lastly, leveraging TSLA, we derive the predicted target-private labels, which are used together with source labels to construct a universal classifier for robust self-training. Compared to methods that rely on the geometric structure of visual representations (Chang et al., 2022; Lai et al., 2023; Wang et al., 2024a), our universal classifier more effectively distinguishes between known and unknown labels and produces more reliable pseudo-labels, leading to further performance enhancements (Sec. 3.6).
To validate our approach, we conducted experiments on four standard UniDA benchmarks, each representing a unique visual recognition domain-adaptation task. In summary, our contributions are:
  1. 1.
    We propose a novel training-free label space alignment method for Universal Domain Adaptation (UniDA), shifting the focus from visual feature alignment to label space alignment using generative vision-language models (VLMs).
  2. 2.
    To address semantic challenges in naïve label space alignment with generative VLMs, we propose new filtering techniques to resolve semantic ambiguities and noise, yielding a refined shared label space between source and target domains, as well as a well-defined target-private label space.
  3. 3.
    By applying the filtered target private classes, we enhance the CLIP zero-shot classifier for UniDA, referred to as the universal classifier, which can be further enhanced through self-training to achieve full adaptation to the target domain.
  4. 4.
    We perform comprehensive analyses to validate the effectiveness of our approach, attaining state-of-the-art performance on H-score and H3-score metrics.
Furthermore, as demonstrated in prior studies, fine-tuning the backbone CLIP on source data diminishes its ability to detect target private classes (Deng and Jia, 2023). This issue, known as feature distortion (Kumar et al., 2022; Shu et al., 2022; Trivedi et al., 2023; Saito et al., 2023), arises when the source data is used for tuning. Our method circumvents tuning on the source data by leveraging the source label space to filter and identify target private labels from the generative VLM. Furthermore, our approach prevents feature distortion, enabling faster experimentation and improved performance.
Figure 3: Comparison between prior UniDA methods and our TLSA+ST. Left: Visual space alignment approaches (Li et al., 2021a; Chang et al., 2022; Lai et al., 2023) rely on cluster matching, which becomes ambiguous under domain shift. Right: Our TLSA+ST discovers labels from target images and aligns them with source labels with adaptive thresholding using the distance in the embedding space between source labels. Since TLSA+ST uses pretrained CLIP’s joint embedding space, it is robust against domain shift.
2 Related Work
2.1 Vision-Language Models
Vision-language models have recently demonstrated great potential in learning generic visual representations. (Radford et al., 2021; Li et al., 2021b; Zhou et al., 2022). Their transferability is attributed to the ability to inject rich information into the model by aligning image-text pairs, rather than just a single word label. For example, CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) demonstrate strong zero-shot performance in visual recognition. Not surprisingly, many studies have achieved high performance within only zero-shot pipelines by using the Vision-Language Models (VLMs) (Adila et al., 2023; Ge et al., 2023; Qu et al., 2024a; Zhao et al., 2024; Zanella and Ben Ayed, 2024; Wang et al., 2024b). Generative VLMs, including BLIP (Li et al., 2022, 2023), GPT-4 (Achiam et al., 2023), and LLaVA (Liu et al., 2024), have demonstrated strong performance on a wide range of vision–language generation tasks. In light of the latest advancements, we utilize BLIP to discover a label agnostic to the source domain for unlabeled target domain samples.
2.2 Universal Domain Adaptation
Unlike UDA, UniDA assumes a distinct scenario where the label set of the target domain often partially intersects with that of the source domain. Determining whether a given data belongs to a known or unknown class is a highly challenging task because the model can only predict the classes it is trained on. Therefore, baseline methods for UDA, such as adversarial learning (Ganin et al., 2016; Tzeng et al., 2017; Long et al., 2018) and metric learning (Long et al., 2015, 2017), lead to degradation in performance. The problem occurs because domain alignment is attempted without considering the possibility that the label space might differ, leading to negative transfer.
With the recent progress of VLMs, there has been growing recognition of the relatively limited exploration compared to pretrained backbones on ImageNet (Russakovsky et al., 2015), such as ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2020). In response, CLIP-Distillation (Deng and Jia, 2023) was proposed, utilizing foundation models (Radford et al., 2021; Oquab et al., 2023) trained on large datasets through self-supervised learning. This approach achieved state-of-the-art performance among CLIP-pretrained backbones by performing auto-calibration via learning the classifier temperature. Similarly, MemSPM (Lai et al., 2023) adopts a memory-based approach that models sub-classes by storing and updating sub-prototypes, while keeping the CLIP backbone frozen. Here, CLIP is employed not for its zero-shot capability, but merely as a generic feature extractor. Fig. 3 illustrates the key differences between prior works and our approach. Unlike previous methods, and as motivated by Fig. 2, we reformulate UniDA as the simpler task of discovering target-private class names. Specifically, we generate candidate labels directly from target-domain samples using a generative VLM, and then apply a filtering strategy that exploits image–text similarity to refine them.
2.3 Out-Of-Distribution Detection
Out-of-Distribution (OOD) detection aims to identify samples during testing that come from distributions not encountered in training. Many research (Ming et al., 2022; Wang et al., 2023) leveraged vision-language foundation models, which integrate both visual and textual information to enhance the performance and flexibility of OOD detection. On the other hand, there is a sub-task of OOD detection known as Open-Set Recognition (OSR). OSR addresses the issue of OOD detection at a class level and is even more closely related to UniDA. Scheirer et al. (2012) first articulated the concept of OSR, which has since inspired extensive research across various fields (Bendale and Boult, 2016; Ge et al., 2017; Oza and Patel, 2019; Vaze et al., 2022; Esmaeilpour et al., 2022; Cen et al., 2023).
Among these, the recent study aims to leverage CLIP’s zero-shot capability for open-set domain adaptation (Zara et al., 2023). However, there are significant limitations when applying this method directly to UniDA. The number of clusters, $k$, is determined based on prior knowledge from the open-set, where the number of clusters exceeds the number of source labels. In a universal set, the decision of $k$ is more challenging as its extent is not pre-defined. Specifically, when using K-means to predict the number of classes, the elbow method is typically applied (Hartigan and Wong, 1979). However, the elbow point is often unreliable due to a smooth score plot, leading to inaccuracies. Furthermore, we argue that texts discovered from the source domain are ineffective for detecting target private classes due to domain shifts, as the shapes of objects can differ entirely between the source and target domains, especially for abstract labels.
Unlike previous methods, our approach leverages instance-level discovery by comparing each discovered label with source labels based on image-text similarity. This eliminates the need to predefine the number of clusters, which would otherwise depend on the specific configurations of datasets. Also, we propose discovering target private classes by examining the relationship between an unbiased text representation and only the target image representation.
2.4 Self-Training
Self-training is a widely studied approach in semi-supervised learning that has recently been explored for enhancing the performance of minority classes (Amini et al., 2024). In situations where domain shift exists, research has primarily focused on Source-free domain adaptation (SFDA) and test-time adaptation (TTA). One of the most well-known approaches is a pseudo labeling technique, which is both simple to implement and intuitive. Pseudo labeling assumes that predictions for confident samples are trustworthy. Initially, training leverages these trustworthy predictions, and as the model advances, the scope of pseudo labeling is progressively broadened. The most important aspect is to maintain class balance. Class imbalance can lead to overconfidence in certain classes, which ultimately degrades the model’s performance. To mitigate this issue, various source-free domain adaptation methods, such as SHOT (Liang et al., 2020) and DIFO (Tang et al., 2024), incorporate a diversity term to perform entropy maximization. ST-SGG (Kim et al., 2024) devises a method to adapt class-adaptive thresholds based on frequency, adjusting them according to the situation in previous iterations.
Inspired by this research direction, we choose to select the top-k most confident samples for each predicted class as pseudo labels to ensure class balance. Furthermore, to maintain this balance among the pseudo labels, we define $k$ as the batch size divided by the number of predicted classes. This approach limits the selection to samples that can be easily classified. Additionally, we empirically confirm that our universal classifier, created by predicting target private labels, discriminates better than the $|C_{s}|+1$ way classifier, where $|C_{s}|$ is the number of source classes (see Fig. 8).
3 Methodology
3.1 Problem Definition
In UniDA, we are given a labeled source domain $\mathcal{D}^{s}=\{(x^{s}_{i},y^{s}_{i})\}^{N_{s}}_{i=1}$ where $x^{s}_{i}\in\mathcal{X}^{s},y^{s}_{i}\in\mathcal{Y}^{s}\subset\mathbb{R}^{|C_{s}|}$, and an unlabeled target domain $\mathcal{D}^{t}=\{x^{t}_{j}\}^{N_{t}}_{j=1}$ where $x^{t}_{j}\in\mathcal{X}^{t}$. We denote the ground-truth labels of the target domain as $\mathcal{Y}^{t}\subset\mathbb{R}^{|C_{t}|}$. $|C_{s}|$ represents the number of source labels, and $|C_{t}|$ represents the number of target labels. $C=C_{s}\cap C_{t}$ denotes the common label sets shared by source and target domains. $\bar{C_{s}}=C_{s}-C$ and $\bar{C_{t}}=C_{t}-C$ stand for source-private label and target private label, respectively. Our goal is to filter the raw discovered label $R$ from the generative VLM to align the ground-truth target private classes $\bar{C_{t}}$.
Figure 4: Overview of TLSA. We first gather candidate target-private labels from a generative VLM, then refine them to train a universal classifier. (a) Synonym label alignment (Sec. 3.3) removes source label synonyms via WordNet: (b) Semantic label alignment (Sec. 3.4) re-scores image-label pairs in the embedding space and decides shared vs. target-private; (c) Frequency-based noisy candidate filtering (Sec. 3.5) prunes noisy low support candidates using frequency banks; and (d) Self-training (Sec. 3.6) applies a teacher-student scheme on class-balanced top-k confident samples.
3.2 Overview
We verified that selecting target private labels derived from discovered labels substantially boosted performance under domain and label space shifts, as displayed in Fig. 2. We introduce TLSA, which iteratively aligns ambiguously similar labels with source labels from the discovered labels and ultimately filters them based on their occurrence patterns after all iterations. Following the TLSA pipeline, we employ a self-training strategy to further enhance discrimination and performance. The overall framework is depicted in Fig. 4.
$\bullet$ Step 1. The goal of this step is to filter the discovered labels by explicitly identifying and removing lexical synonyms of source labels using the WordNet database (Sec. 3.3 and Fig. 4a).
$\bullet$ Step 2. The objective of this step is to remove discovered labels that are semantically ambiguous or resemble the source labels by identifying a similar label set in CLIP’s joint embedding space (Sec. 3.4 and Fig. 4b).
We iterate Step 1 and Step 2 in each cycle until the entire domain is comprehensively explored. We define the refined label set as target private candidate labels.
$\bullet$ Step 3. To address rare instances missed in the previous phase, we removed labels based on their occurrence frequency, operating under the assumption that infrequent target private candidate labels were likely incorrect target private labels (Sec. 3.5 and Fig. 4c).
$\bullet$ Step 4. This step aims to fully adapt the zero-shot model to the target domain, thereby enhancing its performance beyond zero-shot inference. This enhancement is achieved by utilizing a universal classifier weighted by CLIP text embeddings (Sec. 3.6 and Fig. 4d).
3.3 Synonym Label Alignment
In this subsection, we introduce a method for generating raw labels and eliminating synonyms. To acquire the discovered labels, we employ a generative VLM to generate labels for unknown target images. Furthermore, we use WordNet (Miller et al., 1990) to remove semantically ambiguous labels (i.e., synonyms) from the discovered labels, ensuring their alignment with the source label space, as illustrated in Fig. 4a.
Discovering Labels on Unlabeled Target Images. To ensure reliable labels for the input images, we query each target sample $x^{t}_{i}$ using the BLIP-VQA model (Li et al., 2022, 2023). Specifically, we utilize $k=5$ prompts that are semantically equivalent yet lexically diverse, as presented in Tab. 14. The design of these prompts is inspired by the principles of prompt engineering in CLIP (Radford et al., 2021), employing multiple textual variants to reduce prompt sensitivity. Consider prompts like “What is the exact name of the object in the painting?” and “Identify the object in this photo with its exact name,” which share identical semantics but vary in phrasing. The final label was determined via majority voting over $k$ responses. Responses that are overly specific (e.g., specifying a color) or do not correspond to a single valid label are filtered out using a rule-based algorithm. Although a comprehensive sensitivity analysis across different $k$ values was not performed, preliminary observations suggested that $k=5$ struck a balance between computational efficiency and stability; hence, we consistently employed this setting throughout our experiments.
Synonym Label Alignment using WordNet. WordNet (Miller et al., 1990) is a widely utilized lexical database that provides taxonomic relationships among words. Path-based similarity measures are computed by assessing the distances between two synsets within the hierarchy. Consequently, we employed WordNet as an essential filtering step to eliminate noisy labels in $R$.
3.4 Semantic Label Alignment
This subsection elaborates on eliminating semantically ambiguous labels, such as hypernyms, hyponyms, and non-WordNet synonyms. We employed semantic similarity to evaluate the relationship between the text representations of the source and candidate labels, based on their embedding distances from the input images. If candidate labels have a similar proximity to the image as the source labels, they are deemed identical to the corresponding source label, as displayed in Fig. 4b.
[…]
$\mathbf{S}=\begin{pmatrix}s_{1,1}&\cdots&s_{1,|C_{s}|}&s_{1,|C_{s}|+1}&\cdots&s_{1,|\tilde{R}|}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ s_{N,1}&\cdots&s_{N,|C_{s}|}&s_{N,|C_{s}|+1}&\cdots&s_{N,|\tilde{R}|}\end{pmatrix}$ (1)
where $N$ denotes the batch size. The left portion of $\mathbf{S}$ reflects the cosine similarity between the embedding of the images and source labels, while the right portion represents the cosine similarity between images and the filtered discovered labels.
Figure 5: Illustration of semantic label alignment. In this step, we establish the relationship between discovered labels and source label predictions within the prediction set $\mathcal{C}$, as defined by Eq. 5, for a given image $x_{i}$. (a) When source labels are present in $\mathcal{C}$, the input image is classified as belonging to one of the source classes. (b) When no source labels exist in $\mathcal{C}$, the input image is categorized as a target private class. Subsequently, we maintain a frequency count of classes in the frequency bank to enable noisy candidate filtering.
Identification of Target Private Labels from Filtered Discovered Labels. Using semantic similarity, we aim to determine whether an input image belongs to the source or target private classes. An overview of this procedure is presented in Fig. 5. When determining whether a label originates from a source-private or target, we found that naive alignment methods—such as top-1 prediction-based alignment and margin-based alignment—performed poorly in our experiments (see Tab. 8). For instance, when ‘backpack’ is the source label and ‘bag’ is the discovered label (Fig. 2c), ‘bag’ and ‘backpack’ are not synonyms in WordNet; however, within the Office31 context, they effectively represent the same class. Treating them as distinct categories leads to numerous ‘backpack’ samples being misclassified as ‘bag’, resulting in artificial label splits. In such cases, top-1 predictions alternate between ‘backpack’ and ‘bag’ due to uncontrollable image conditions, without signifying any meaningful semantic difference. These semantically ambiguous labels might be included among certain target private labels. To address this, the margin value between the top-1 and top-2 predicted probabilities can evaluate the distance between source labels and the filtered discovered labels. However, this method relies on a predefined fixed threshold to assess reliability, ignoring class similarities that often vary adaptively across datasets and domains.
To overcome this limitation, we adaptively align the filtered discovered labels with source labels, removing the need for a fixed threshold. By analyzing the similarity matrix $S$, we first eliminate noisy labels and identify only reliable predictions within the top $k$ predictions. We define a prediction set $\mathcal{C}$ as the collection of refined and filtered predictions from the top $k$ predictions using the two proposed criteria.
$\bullet$ Criterion 1: Maximum-gap-based thresholding. A prominent gap in similarity scores was identified among the top $k$ rankings. The labels preceding this gap accurately correspond to multiple image predictions. Utilizing this criterion, we establish the meaningful prediction set $\mathcal{C}$. The threshold is defined as follows:
$J=\underset{j\in\{1,\dots,k-1\}}{\arg\max}\left(s_{i,j-1}-s_{i,j}\right),$ (2)
$\tau^{i}_{\text{gap}}=s_{i,J},$ (3)
where i represents the image index, j the ranking position, and $J$ denotes the point at which a significant gap occurs.
$\bullet$ Criterion 2: Average score-based thresholding. We define the prediction set as comprising scores that exceed the average score of the top $k$ predictions. The threshold is defined as the average of the top $k$ scores:
$\tau^{i}_{avg}=\frac{\sum_{j=1}^{k}{s_{i,j}}}{k}.$ (4)
Empirically, we maximize the prediction set size by selecting the lower of the two criteria (i.e., $\tau^{i}_{gap}$ and $\tau^{i}_{avg}$). This approach enhances the inclusion of source labels within the label set, thereby reducing the likelihood of ambiguous labels being identified as private target candidates. In other words, by selecting the smaller threshold between $\tau_{\text{gap}}$ and $\tau_{\text{avg}}$, the criterion for classifying a sample as target private becomes more stringent. The threshold $\tau_{\text{set}}$ and prediction set $\mathcal{C}$ for each instance are formulated as follows:
$\tau^{i}_{set}=\min(\tau^{i}_{gap},\tau^{i}_{avg}),$ (5)
$\mathcal{C}^{i}=\{\mathbf{t}_{i,j}|j\in I^{i}_{k}\land s_{i,j}>\tau_{set}^{i}\},$ (6)
where $\mathbf{t}$ represents the text comprising extensions of source labels and filtered discovered labels, and $I^{i}_{k}$ denotes the top-$k$ indices for the $i^{\text{th}}$ sample. Utilizing the selected prediction set $\mathcal{C}$, we perform a source label presence check to achieve semantic label alignment. Specifically, if a source label is present in $\mathcal{C}^{i}$ (Fig. 5a), sample $i$ is classified as a source sample. When it is absent from the prediction set $\mathcal{C}$, sample $i$ is identified as a target private sample (Fig. 5b).
3.5 Frequency-based Noisy Candidate Filtering
Noisy and incorrectly predicted private target candidate labels may still persist. To address this, we developed a method that filters these noisy labels by analyzing their frequencies, as shown in Fig. 4c. We hypothesized that low-frequency classes are likely noisy labels in the target domain. To precisely measure label frequency, we employed a frequency bank. For samples classified as source classes, the top-1 prediction $p_{i}$ is stored in the frequency bank as a reference standard (Fig. 5a). Similarly, for samples classified as target private classes, the discovered label $r_{i}$ is added to the frequency bank even if $r_{i}\notin\mathcal{C}^{i}$(Fig. 5b). The rationale for storing $r_{i}$, the label identified by BLIP, instead of $p_{i}$, the closest label in the CLIP space, is detailed in Appendix D. The frequency bank is defined as $F=\{c:f_{c}\}$, where $f_{c}$ represents the frequency of class $c$ after semantic label alignment, including both source and filtered discovered labels. Considering the presence of noisy labels in the frequency bank, we used the count of the most frequent class rather than the median or mean of the label frequency distribution. We set a threshold $\tau_{\text{freq}}$ as the fixed number of samples in the most frequent class. This scaling factor, denoted by $\epsilon$, is a dataset-agnostic parameter. Finally, the predicted target private labels $\mathcal{C}_{p}$ are derived from the raw discovered labels $R$ as follows:
$\tau_{freq}=\epsilon\times\max(F),$ (7)
$C_{p}=\{c|f_{c}>\tau_{freq}\land c\notin C_{s}\}.$ (8)
We employ an empirically determined value of $\epsilon=0.01$ across all datasets, as presented in Tab. 10.
By integrating complementary components of the pipeline—lexical filtering validated through WordNet, semantic filtering within a joint image–text embedding space, and a frequency-based filtering step that removes residual noisy discovered labels depending on dataset scale—our method achieves strong complementarity among the three, as evidenced in Tab. 7.
3.6 Self-Training with Universal Classifier
In UniDA, the prior typically uses randomly initialized prototype-based classifiers (Li et al., 2021a; Chang et al., 2022; Lai et al., 2023). However, this method complicates determining the number of target clusters, which corresponds to the number of private target classes. In contrast, we utilize a universal classifier that automatically determines the number of target private classes, as depicted in Fig. 4d. Furthermore, we applied self-training, a prevalent technique in source-free domain adaptation (Liang et al., 2020; Zhang et al., 2023; Karim et al., 2023; Tang et al., 2024). This method reduces bias toward the source domain, offering an advantage over approaches that depend on visual-space domain alignment and source-supervised learning. Additionally, we selected reliable pseudo labels based on the top-$k$ confidence samples per class and employed a student-teacher framework with EMA updates.
Creating a Universal Classifier. Our inference leverages the weights of the universal classifier, denoted by $\mathbf{z}_{C_{s}}$ and $\mathbf{z}_{C_{p}}$, which are produced by inputting $C_{p}$ into the text encoder of CLIP. For post-processing, instances classified as $C_{s}$ were retained unchanged, while those labeled as $C_{p}$ were adjusted to $|C_{s}|+1$, ensuring a threshold-free method.
Reliable Pseudo Label Selection. To leverage pseudo-labels, we trained a student model to enhance the teacher model’s performance using delayed updates. However, naive pseudo-labeling tends to select only easily classifiable classes, leading to overfitting of these simpler categories. To address this, we introduce a balanced pseudo-label selection strategy that improves the generalization of the predictions of the pretrained model. Rather than requiring a large dataset, our approach prioritizes maximum diversity by selecting the top-k (%) most confident pseudo-labels per class, as proposed by Zara et al. (2023). Nonetheless, even with a top-k (%) threshold, more prevalent classes are disproportionately selected, which can degrade performance. To mitigate this, we define a balanced $k$ value to select only the top-$k$ most confident samples for each class. Specifically, $k$ is calculated as $k=\text{min}\{n_{c_{i}},N/n_{c}\}$, where $n_{c}$ is the number of unique predicted classes and $n_{c_{i}}$ is the number of predictions for class $c_{i}$.
Robust Self-Training. To ensure robust self-training, we apply Exponential Moving Average (EMA) updates to the teacher model, defined by $\theta_{t}=(1-\alpha)\theta_{s}+\alpha\theta_{t}$. Here, $\theta_{t}$ and $\theta_{s}$ represent the parameters of the teacher and student models, respectively, effectively mitigating confirmation bias. The student model updates with each iteration, whereas the teacher model updates once per epoch, as defined by the preceding equation. This configuration is crucial for the performance of the model (Sec. 5). Additionally, we implemented adapter tuning to limit further tuning capabilities. We then perform self-training using the following loss function: Let $M$ represent the set of samples employed as pseudo-labels, $\hat{y}^{T}_{t,i}$denotes the output of the teacher model, and $\hat{y}^{S}_{t,i}$represents the output of the student model.
$\mathcal{L}_{self}=-\sum_{i\in M}{\hat{y}^{T}_{t,i}\log{\hat{y}^{S}_{t,i}}},$ (9)
By utilizing a frozen classifier, a common approach in SFDA (Liang et al., 2020), adapter tuning mitigates error accumulation by addressing the delayed conversion rate.
4 Experiments
This section outlines the dataset, implementation, experimental settings, and results.
Datasets. We utilize four benchmark datasets for visual recognition domain adaptation: Office31 (O), Office-Home (OH), VisDA (VD), and DomainNet (DN). Office31 (Saenko et al., 2010) is a small-scale benchmark dataset comprising images from 31 categories distributed across three domains: Amazon (2,817 images), Dslr (498 images), and Webcam (795 images). Office-Home (Venkateswara et al., 2017) is a benchmark dataset designed for domain adaptation, featuring images from 65 categories across four domains: Art (9,655 images), Clipart (9,516 images), Product (15,688 images), and Realworld (14,840 images). VisDA (Peng et al., 2017) is a benchmark dataset for domain adaptation, containing images from 12 labels distributed across two domains: Synthetic (152,397 images) and Real (55,088 images). DomainNet (Peng et al., 2019) is a large-scale benchmark dataset for domain adaptation, notable for its extensive size. It includes images from 345 labels distributed across six domains: Clipart, Infograph, Painting, QuickDraw, Real, and Sketch. The sheer scale of DomainNet makes it a particularly challenging and comprehensive dataset for evaluating domain adaptation methods. We conducted experiments using only three domains from DomainNet that are commonly employed in existing methods for UniDA - Painting, Sketch, and Real.
Table 1: Split-setting of our experiments. Details of experimental split settings indicating the counts of source-private and shared classes (Source-Private / Shared) in each dataset under different adaptation scenarios (open-partial, open, closed, partial).
Datasets Split settings
open-partial open closed partial
Office31 10/10 0/10 0/31 21/10
Office-Home 5/10 0/15 0/65 40/25
VisDA 3/6 0/6 0/12 6/6
DomainNet 50/150 0/150 0/345 195/150
[…]
Implementation Details. We use CLIP (Radford et al., 2021) backbone with ViT-B/16 (Dosovitskiy et al., 2020) as our image encoder. For the generative VLM, we utilize BLIP (Li et al., 2022), a model pretrained on VQAv2 (Goyal et al., 2017) within the LAVIS library [^1] https://github.com/salesforce/LAVIS. This model incorporates a visual encoder based on ViT-Base (Dosovitskiy et al., 2020) and employs a text encoder and decoder that adhere to the BERT (Devlin et al., 2018) architecture. Most of the baseline code and pipelines are based on the uniood repository [^2] https://github.com/szubing/uniood.
Our experimental class split settings follow the standard protocol used by many existing UniDA approaches (Saito et al., 2020; Saito and Saenko, 2021; Chang et al., 2022; Lu et al., 2024; Qu et al., 2024b, c; Deng and Jia, 2023), as detailed in Tab. 1.
During the alignment phase, to expedite filtering, we conducted experiments with a batch size of 128. These experiments were executed on a single H100 GPU, consuming approximately 15GB of memory during the filtering process. As the source domain was not trained separately, the overall experimental time was significantly reduced.
In the self-training phase, we conducted experiments with a batch size of 64, a learning rate of 0.01, and an adapter with bottleneck dimensions of r=64, adhering to the default settings of AdaptFormer (Chen et al., 2022b). All baseline methods were implemented with identical iterations, batch sizes, and learning rates using a frozen backbone, as described in Deng and Jia (2023). The iterations were set to 2,500; 5,000; and 10,000, depending on the task setting, as presented in Tab. 2, following (Deng and Jia, 2023). Our methods required approximately 30 minutes per domain, allowing us to complete all experiments on the entire dataset within a day.
Table 2: Training iterations on different split settings. We follow the same setting in (Deng and Jia, 2023). Since we use a batch size 64, which is twice the original value, we reduce the number of iterations accordingly by half.
Datasets Split settings
open-partial open closed partial
Office31 2500 2500 5000 5000
Office-Home 2500 2500 5000 5000
VisDA 5000 5000 10000 10000
DomainNet 5000 5000 10000 10000
Evaluation Metric. We use the existing two evaluation metrics for UniDA in addition to classification accuracy. H-score (Fu et al., 2020) is the harmonic mean of the instance accuracy on common class $a_{C}$ and accuracy on the single private class $a_{\bar{C_{t}}}$. This metric effectively captures the trade-off between known-class accuracy and out-of-class detection accuracy. Moreover, we also use the H3-score. The H3-score is first proposed by UniOT (Chang et al., 2022) and extends the H-Score by incorporating the Normalized Mutual Information (NMI), a well-known measure that indicates how well the target-private clusters are formed. In the PDA and CDA settings where target-private instances do not exist, the average class accuracy is used in place of the H-score. Similarly, the $H^{3}$ score is also substituted with this measure.
5 Results
Table 3: H-score (%) and H3-score (%) comparisons between existing methods and the proposed method using the CLIP ViT-B/16 backbone across four UniDA settings (open-partial, open, closed, partial). An asterisk (*) represents results for some methods referenced from the previous paper (Lai et al., 2023). A dash (-) represents that any metrics not reported in the official papers.
Methods Office31 Office-Home VisDA DomainNet Avg (10/10) (10/0) (10/21) (31/0) (10/5) (15/0) (25/40) (65/0) (6/3) (6/0) (6/6) (12/0) (150/50) (150/0) (150/195) (345/0) H-score SO 85.8 86.8 83.8 70.8 79.1 76.5 51.8 49.6 67.8 71.4 49.1 51.1 55.8 59.7 32.9 34.8 62.9 DANCE 91.7 93.2 73.8 67.7 83.6 78.3 46.9 49.4 66.6 76.6 33.7 46.3 55.1 59.4 30.0 34.1 61.6 OVANet 89.0 89.3 84.0 65.9 80.5 76.3 58.8 56.1 56.6 44.4 36.6 40.9 62.0 64.4 45.0 46.2 62.2 UniOT 82.3 93.9 31.5 54.2 86.3 81.8 31.8 46.4 58.2 77.3 33.6 54.2 65.6 69.4 40.9 51.5 59.8 WiSE-FT 70.4 88.3 38.5 30.9 72.8 65.8 11.7 8.8 54.7 67.8 19.1 22.2 2.4 6.0 0.19 0.2 35.0 CLIP-Distill 83.5 84.3 85.6 73.7 84.7 81.3 74.4 69.0 82.8 82.2 78.7 71.1 64.8 66.7 56.4 51.3 74.4 GLC* 84.8 90.1 91.5 - 82.6 86.2 75.4 - 80.3 81.6 86.2 - 46.8 - - - - MemSPM* 88.5 95.6 94.4 - 84.2 87.9 74.5 - 79.3 79.7 87.9 - 56.7 - - - - TLSA 89.9 92.6 96.1 80.5 91.3 89.2 82.0 79.0 86.3 83.9 81.8 72.6 75.4 76.9 70.1 68.9 82.3 TLSA+ST 91.9 95.2 96.9 82.9 93.5 91.4 85.3 81.8 87.6 86.0 83.0 75.6 74.3 74.9 74.4 67.9 83.9 H3-score SO 83.4 85.8 83.8 70.8 77.2 75.2 51.8 49.6 72.6 73.4 49.1 51.1 59.6 62.5 32.9 34.8 63.3 DANCE 87.0 89.8 73.8 67.7 80.5 77.7 46.9 49.4 71.7 76.9 33.7 46.3 59.1 62.2 30.0 34.1 61.6 OVANet 85.2 87.3 84.0 65.9 78.0 75.2 58.8 56.1 63.6 51.8 36.6 40.9 64.1 65.8 45.0 46.2 62.8 UniOT 80.7 89.8 31.5 54.2 82.3 79.7 31.8 46.4 61.3 75.8 34.0 53.6 66.6 69.0 40.9 51.5 59.3 WiSE-FT 72.5 86.6 38.5 30.9 72.0 68.1 11.7 8.8 71.5 78.3 19.1 22.2 3.6 8.6 0.2 0.2 37.1 CLIP-Distill 81.7 83.8 85.6 73.7 80.2 78.3 74.4 69.0 81.8 86.0 78.7 71.1 66.1 67.4 56.4 51.3 74.1 TLSA 85.8 89.1 96.1 80.5 84.8 82.9 82.0 79.0 85.7 81.7 81.8 72.6 73.0 73.9 69.6 68.2 80.5 TLSA+ST 85.3 89.9 96.9 82.9 90.2 89.0 85.3 81.8 88.4 85.0 83.0 75.6 75.4 75.5 74.4 67.9 83.2
Table 4: H-score (%) comparison in OPDA scenarios on the Office-Home dataset. Results for some methods are referenced from GLC++ (Qu et al., 2024c) and LEAD (Qu et al., 2024b). Here, SF indicates source data-free methods, and SM denotes approaches that use a source model. We compare TLSA with both source-free methods and those using source models. An asterisk (*) indicates reproduced results.
Backbone Methods SF SM Office-Home Avg
Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr
ImageNet SO 47.3 71.6 81.9 51.5 57.2 69.4 56.0 40.3 76.6 61.4 44.2 73.5 60.9
UANh - 51.6 51.7 54.3 61.7 57.6 61.9 50.4 47.6 61.5 62.9 52.6 65.2 56.6
CMUi - 56.0 56.9 59.2 67.0 64.3 67.8 54.7 51.1 66.4 68.2 57.9 69.7 61.6
DANCEa - 61.0 60.4 64.9 65.7 58.8 61.8 73.1 61.2 66.6 67.7 62.4 63.7 63.9
DCCj - 58.0 54.1 58.0 74.6 70.6 77.5 64.3 73.6 74.9 81.0 75.1 80.4 70.2
GATEk - 63.8 75.9 81.4 74.0 72.1 79.8 74.7 70.3 82.7 79.1 71.5 81.7 75.6
OVANetb - 61.7 77.3 80.6 68.9 68.7 61.9 50.4 47.6 61.5 62.9 52.6 65.2 56.6
UniOTc - 67.3 80.5 86.0 73.5 77.3 84.3 75.5 63.3 86.0 77.8 65.4 81.9 76.6
GLCf 64.3 78.2 89.8 63.1 81.7 89.1 77.6 54.2 88.9 80.7 54.2 85.9 75.6
MLNetl - 68.2 83.8 85.0 73.6 78.2 82.2 75.2 64.7 85.1 78.8 69.9 83.9 77.4
LEADm - 62.7 78.1 86.4 70.6 76.3 83.4 75.3 60.6 86.2 75.4 60.7 83.7 75.0
CLIP SO* 69.8 80.3 83.8 80.5 81.5 85.3 77.5 71.2 84.7 80.5 76.0 77.6 79.1
DANCE*a - 78.0 86.4 87.1 81.1 85.2 85.6 82.4 79.9 90.5 83.6 79.6 84.1 83.6
OVANet*b - 72.7 88.0 86.4 80.6 84.0 85.3 75.2 63.6 89.1 82.0 74.4 85.0 80.5
UniOT*c - 85.5 89.8 91.1 79.9 87.3 89.3 80.8 83.5 88.9 83.5 84.3 91.6 86.3
WiSE-FT*d 59.4 86.9 87.0 57.6 80.3 80.9 62.7 50.2 85.1 71.0 64.7 88.4 72.8
GLCf 79.4 88.9 90.8 76.3 84.7 89.0 71.5 72.9 85.7 78.2 79.4 90.0 82.6
DCCj - 62.6 88.7 87.4 63.3 68.5 79.3 67.9 63.8 82.4 70.7 69.8 87.5 74.4
MemSPMg - 78.1 90.3 90.7 81.9 90.5 88.3 79.2 77.4 87.8 78.8 76.2 91.6 84.2
CLIP-Distill*e 81.4 79.6 86.1 87.9 80.3 86.7 87.0 81.3 80.6 85.2 82.0 88.9 84.3
TLSA 87.4 95.2 95.8 86.8 95.2 95.8 86.8 87.4 95.8 86.8 87.4 95.2 91.3
TLSA+ST 90.2 96.3 96.7 90.6 96.3 96.7 90.6 90.2 96.7 90.6 90.2 96.3 93.4
Table 5: H-score (%) comparison in OPDA scenarios on the Office31, DomainNet, and VisDA datasets, respectively. MLNet (Lu et al., 2024) does not provide experimental results for DomainNet, any metrics not reported in their paper are marked as “-”. An asterisk (*) indicates reproduced results.
Backbone Methods SF SM Office31 DomainNet VisDA
A2D A2W D2A D2W W2A W2D Avg P2R P2S R2P R2S S2P S2R Avg S2R
ImageNet SO 70.9 63.2 39.6 77.3 52.2 86.4 64.9 57.3 38.2 47.8 38.4 32.2 48.2 43.7 25.7
UANh - 59.7 58.6 60.1 70.6 60.3 71.4 63.5 41.9 39.1 43.6 38.7 38.9 43.7 41.0 34.8
CMUi - 68.1 67.3 71.4 79.3 72.2 80.4 73.1 50.8 45.1 52.2 45.6 44.8 51.0 48.3 32.9
DANCEa - 78.6 71.5 79.9 91.4 72.2 87.9 80.3 21.0 37.0 47.3 46.7 27.7 21.0 33.5 42.8
DCCj - 88.5 78.5 70.2 79.3 75.9 88.6 80.2 56.9 43.7 50.3 43.3 44.9 56.2 49.2 43.0
GATEk - 87.7 81.6 84.2 94.8 83.4 94.1 87.6 57.4 48.7 52.8 47.6 49.5 56.3 52.1 56.4
OVANetb - 85.8 79.4 80.1 95.4 84.0 94.3 86.5 56.0 47.1 51.7 44.9 47.4 57.2 50.7 53.1
UniOTc - 83.7 85.3 71.4 91.2 70.9 90.8 82.2 59.3 47.8 51.8 46.8 48.3 58.3 52.0 57.3
GLCf 81.5 84.5 89.8 90.4 88.4 92.3 87.8 63.3 50.5 54.9 50.9 49.6 61.3 55.1 73.1
MLNetl - 83.9 78.5 81.3 95.4 83.2 96.4 86.5 - - - - - - - 69.9
LEADm 85.4 85.0 86.3 90.9 86.2 93.1 87.8 59.9 46.1 51.3 45.0 45.9 56.3 50.8 76.6
CLIP SO* 84.4 79.0 88.1 93.3 80.2 89.8 85.8 64.2 50.2 53.1 52.1 48.3 67.1 55.8 67.8
DANCE*a - 91.6 84.7 93.7 95.8 87.8 96.4 91.7 64.2 50.1 50.8 51.2 47.4 67.2 55.1 66.6
OVANet*b - 86.8 76.1 90.1 94.9 91.0 95.2 89.0 68.2 58.6 60.5 59.1 57.0 68.6 62.0 56.6
UniOT*c - 78.5 73.4 82.5 88.4 83.0 88.1 82.3 71.4 62.8 62.5 64.6 59.5 72.7 65.6 56.3
WiSE-FT*d 60.4 50.7 72.6 77.9 70.6 90.2 70.4 5.07 1.15 2.84 1.64 0.84 3.06 2.43 46.5
DCCj - 82.2 76.9 83.6 75.2 85.8 88.7 82.1 61.1 38.8 51.8 49.3 49.1 60.3 52.2 61.2
GLCf 80.5 80.4 77.5 95.6 77.7 96.9 84.8 51.2 44.5 55.6 43.1 47.0 39.1 46.8 80.3
MemSPMg - 88.0 84.6 88.7 87.6 87.9 94.3 88.5 62.4 52.8 58.5 53.3 50.4 62.6 56.7 79.3
CLIP-Distill*e 81.6 81.1 87.7 81.0 87.6 81.8 83.5 74.2 61.6 58.5 60.2 59.7 74.8 64.8 81.9
TLSA 92.7 86.8 90.1 86.8 90.1 92.6 89.9 83.6 71.2 71.3 71.2 71.3 83.6 75.4 86.3
TLSA+ST 95.4 88.7 91.5 88.7 91.5 95.4 91.9 82.3 69.4 71.2 69.4 71.2 82.3 74.3 87.6
5.1 Quantitative Comparison
Comparison with Baseline Methods. We compared TLSA against existing UniDA approaches. Among these, CLIP-Distillation achieves state-of-the-art (SOTA) performance. Except for the Office31 open-set and VisDA partial settings, we attain SOTA across all scenarios. Although the impact of self-training is more significant on certain datasets, it generally enhances the H3-score in most cases. The results are summarized in Tab. 3. Since calculating the H-score and H3-score for partial and closed-set scenarios is infeasible, we employed the average accuracy score as a substitute for the H-score and H3-score, following Deng and Jia (2023). These findings highlight the versatility and effectiveness of TLSA and TLSA+ST (Self-Training) across diverse UniDA settings, particularly in open-partial and open scenarios where managing unseen classes is critical.
In the open-partial setting, TLSA+ST achieved the highest scores across datasets, with H-scores of 93.5 for Office-Home and 87.3 for VisDA. The H3-scores were similarly high, reaching 89.2 on Office-Home and 87.6 on VisDA, demonstrating the ability of TLSA+ST to adapt to partially overlapping classes. In the open setting, TLSA+ST leads with H-scores of 95.2 on Office31 and 87.6 on VisDA, while H3-scores are also competitive at 89.9 on Office31 and 84.9 on VisDA, showcasing the robustness of the method in managing unseen target classes. In the closed setting, where all classes overlap, TLSA+ST maintained competitive performance with an H-score of 75.6 on VisDA, significantly surpassing other methods. In the partial setting, TLSA+ST also delivered strong results, achieving an H-score of 74.4 on DomainNet. Collectively, these results highlight the effectiveness of TLSA and TLSA+ST across various UniDA scenarios, particularly excelling in open-partial and open settings where handling unseen classes is crucial.
Figure 6: Robustness across different class split settings in Office-Home (left) RealWorld to Clipart, and DomainNet (right) Real to Painting.
Figure 7: The number of predicted target-private labels across different split settings and datasets.
Tab. 4 and 5 compare our results with existing methods in the open partial domain adaptation scenario using both ResNet50 (ImageNet-pretrained) and ViT-B/16 (CLIP-pretrained) backbones. Our approach achieved state-of-the-art performance across nearly all domain-to-domain adaptations. However, a slight reduction in H-score was observed due to self-training in DomainNet. While closed-set accuracy improved, OOD detection performance tended to decrease, and vice versa. This indicates a trade-off: self-training shifts the focus of the model, often leading to a decline in H-score performance.
[…]
H-score: 41.7%
(a) CLIP zero-shot
H-score: 66.02%
(b) TLSA+ST ($|C_{s}|{+}1$ classifier)
H-score: 90.2%
(c) TLSA+ST (Universal classifier)
Figure 8: The t-SNE plots of target domain samples for the zero-shot and TLSA+ST methods in the Office-Home RealWorld $\to$ Clipart OPDA setting. Blue points indicate target-private samples, while red points represent source-class samples.
Table 6: List of ground truth target-private labels and final candidates extracted from TLSA for the Office31, DSLR, open-partial setting. The same color indicates the same semantic group.
True target-private Predicted target-private
[Uncaptioned image] pen pen bat
pencil -
[Uncaptioned image] printer printer fax machine
[Uncaptioned image] ruler ruler -
[Uncaptioned image] punchers stapler -
[Uncaptioned image] ring binder fax machine -
[Uncaptioned image] scissors scissors -
[Uncaptioned image] tape dispenser paper dispenser -
[Uncaptioned image] trash can trash can recycle bin
[Uncaptioned image] stapler stapler -
[Uncaptioned image] phone modem -
[Uncaptioned image] speaker speaker usb drive
- tv scale
remote
5.2 Qualitative Comparison
Analysis of the Number of Predicted Target-Private Labels. Fig. 7 displays the number of predicted target-private labels in our study. Across various datasets and settings, our method accurately predicts target-private labels compared to their ground-truth counterparts. This illustrates the flexibility of our design in adapting to different split configurations and datasets.
Feature Visualization. We present t-SNE (Long et al., 2016) plots in Fig. 8 to compare feature discrimination between the CLIP zero-shot and $|C_{s}|+1$classifiers using our approach. As other methods employ a fixed backbone identical to that in (Deng and Jia, 2023), we omitted their representations from the visualization. Furthermore, to emphasize our universal classifier’s discriminative ability, we utilized a $|C_{s}|+1$classifier with LP-ST (linear probing via self-distillation and self-training using top-$k$ confident sample selection), similar to LP-FT (Kumar et al., 2022). The $|C_{s}|+1$-way classifier underperforms compared to TLSA due to its constrained capacity to represent unknown classes.
Predicted Target-Private Labels Analysis. Tab. 6 presents examples of the target-private set extracted by the proposed method. We observed that these predicted labels align with the ground-truth target-private categories in the target domain under the DSLR setting of Office31, indicating that TLSA effectively identifies genuine target-private classes. Importantly, the discovered labels do not need to precisely match the ground-truth annotations; they often also capture prominent attributes of target-domain images. This flexibility is beneficial because generative VLMs can function effectively even on objects not encountered during pretraining, making them particularly suited for open vocabulary or unknown classes. This process generates a set of discovered labels, $R=\{r_{i}\}_{i=1}^{N}$, where $N$ represents the batch size of the target image.
5.3 Ablation Study
Ablation Study and Comparison of the Alignment Pipeline. We implement an ablation study of label space alignment. As can be seen in Tab. 7, Synonym label alignment serves as a fundamental filtering method, contributing to performance to a certain extent.
Top-1 prediction-based alignment relies solely on the top-1 prediction. If the top-1 prediction corresponds to a discovered label, the sample is classified as a target-private sample; otherwise, if it corresponds to a source label, it is identified as a shared sample. In contrast, margin-based alignment requires that, even when a discovered label is the top-1 prediction, it is considered a target-private candidate only if its score exceeds that of the top-1 source label by a specified margin. This margin is empirically fixed at $\tfrac{1}{|C_{s}|}$ across all instances, which yields the best performance for this method.
Frequency-based noisy candidate filtering is a vital component of our pipeline; without it, highly noisy candidate labels could be incorrectly designated as target-private labels. A comparison of these methods is shown in Tab. 7.
Table 7: Ablation study with different components of our alignment pipeline approach. The semantic alignment strategy refers to the comparison of alignment methods used in step 2. Top-1 refers to top-1 prediction-based alignment, Margin refers to margin-based alignment, and TLSA refers to our method.
Step 1 Step 2 Step 3 Office31 Office-Home VisDA DomainNet
Top-1 Margin Ours OPDA ODA PDA CDA OPDA ODA PDA CDA OPDA ODA PDA CDA OPDA ODA PDA CDA
88.8 91.5 89.9 71.3 86.6 83.8 72.7 66.3 76.6 75.2 64.7 54.1 74.8 75.8 66.2 65.3
89.2 92.1 91.6 71.7 87.3 84.5 72.9 66.4 77.8 77.4 67.2 57.2 75.1 76.1 66.6 65.5
88.6 92.0 92.3 76.8 87.2 86.4 76.9 72.2 86.3 82.2 75.6 70.8 76.1 77.1 69.1 67.6
90.0 91.9 94.6 78.6 89.8 85.2 79.7 74.2 47.3 45.7 28.9 28.8 71.6 71.1 62.8 60.4
89.9 92.6 96.1 80.5 91.3 89.2 82.0 79.0 86.3 83.9 81.8 72.6 75.4 76.9 70.1 68.9
Table 8: Ablation study comparing H-scores across different prediction set decision thresholds
methods OPDA ODA PDA CDA Avg
w/o $\tau_{gap}$ 85.8 85.6 81.0 75.0 81.9
w/o $\tau_{mean}$ 84.1 83.7 77.5 71.2 79.1
Ours 85.7 85.6 82.5 75.2 82.3
Table 9: Ablation study with different components of our self-training approach. Delay refers to how long the teacher model’s updates are delayed.
Delay Selection Strategy Office31 Office-Home VisDA DomainNet
Top-k% Balanced Top-k OPDA ODA PDA CDA OPDA ODA PDA CDA OPDA ODA PDA CDA OPDA ODA PDA CDA
Universal Classifier zero-shot 89.9 92.6 96.1 80.5 91.3 89.2 82.0 79.0 86.3 83.9 81.8 72.6 75.4 76.9 70.1 68.9
91.5 94.7 97.2 82.3 93.3 90.9 84.6 81.5 87.1 85.8 85.0 75.7 73.9 74.4 72.1 67.0
91.6 95.1 97.7 82.5 93.7 91.3 85.2 81.2 87.2 86.0 85.4 76.1 73.9 74.4 72.2 66.6
91.5 94.7 97.2 82.3 93.3 90.9 84.6 81.5 87.1 85.8 85.0 75.7 73.9 74.4 72.1 67.0
91.9 95.2 96.9 82.9 93.5 91.4 85.3 81.8 87.6 86.0 83.0 75.6 74.3 74.9 74.4 67.9
Ablation Study of Two Criteria for Label Space Alignment. We conducted an ablation study on the two thresholds, $\tau_{\text{avg}}$and $\tau_{\text{gap}}$, used in adaptive filtering. As outlined in Tab. 8, $\tau_{\text{avg}}$significantly impacts performance more than $\tau_{\text{gap}}$. Furthermore, $\tau_{\text{avg}}$and $\tau_{\text{gap}}$operate as complementary thresholds. Our design ensures robustness against split-setting variations and diverse datasets.
Ablation Study of Self-Training Components. We conducted experiments to enhance the reliability of our self-training design. The key observation was that incorporating delayed EMA components (Delay in Tab. 9) slowed the training process of the teacher model and significantly improved the model performance. Moreover, our approach surpassed the top-k% pseudo-label sampling strategy proposed in a previous study (Zara et al., 2023) (Tab.  9).
5.4 Further Analysis
Table 10: Comparison of H-score on DomainNet to analyze the sensitivity of hyperparameter k in semantic label alignment and $\epsilon$ in frequency-based noisy candidate filtering.
k OPDA ODA PDA CDA Avg
4 75.9 77.3 69.5 68.4 72.7
5 75.4 76.9 70.1 68.9 72.8
6 74.6 76.3 70.4 69.3 72.6
7 72.1 74.8 70.8 69.6 71.8
$\epsilon$ OPDA ODA PDA CDA Avg
0.01 75.4 76.9 70.1 68.9 72.8
0.02 73.4 75.7 70.6 69.6 72.3
0.03 72.1 74.4 71.0 69.7 71.8
0.04 71.1 73.2 71.0 69.8 71.3
This section includes sensitivity analyses of hyperparameters during the alignment process, scalability evaluations relative to backbone size, learning dynamics during self-training, and analyses of trade-offs related to source sample usage.
Sensitivity Analysis for Hyperparameters in TLSA. We analyze the sensitivity of our hyperparameters, $k$, in the top-$k$ prediction within the label space alignment pipeline and $\epsilon$ in frequency-based noisy candidate filtering. The results are presented in Tab. 10. The experiments on the DomainNet dataset determined the optimal value based on average performance.
Specifically, we performed a sensitivity analysis on the $k$ value used for top-$k$ predictions in Step 2 (semantic label alignment) of the filtering pipeline. Our pipeline consistently outperformed other UniDA methods, leading us to select $k=5$ for its robustness across various configurations.
We perform a sensitivity analysis on the value of $\epsilon$ used for frequency-based noisy candidate filtering. A larger $\epsilon$ results in a higher threshold, accepting fewer candidate labels, while a smaller $\epsilon$ leads to a lower threshold, allowing a broader range of target private label candidates. We observe that the performance remains robust across different values of the size of $\epsilon$. Among these, we select 0.01, which shows superior performance across 4 datasets.
Table 11: Comparison of H-score with larger backbones: OH denotes Office-Home, VD denotes VisDA, and DN denotes DomainNet.
Backbone Split Setting
Office31 OH VD DN
(a) TLSA
ViT-B/16 85.7 85.6 82.5 75.2
ViT-L/14@336px 92.9 90.3 84.6 78.9
(b) TLSA+ST
ViT-B/16 86.8 86.9 84.9 77.0
ViT-L/14@336px 94.0 92.0 83.5 79.9
Scalability Analysis for Model Sizes. We conducted experiments using the CLIP-ViT-B/16 backbone. A crucial aspect of validating our method is determining whether performance scales with the backbone size. As illustrated in Tab. 11, increasing the backbone to CLIP-ViT-L/14@336px leads to a significant performance enhancement.
Figure 9: Learning Statistics of Our Self-Training. The x-axis denotes the iteration.
Learning Statistics of Self-Training. We present self-training learning statistics for Office31 DSLR and Office-Home Clipart under the OPDA setting in Fig. 9. A notable performance improvement occurs after a certain number of iterations.
Table 12: Comparison of H-scores between self-training combined with source training and ours (self-training only)
Method OPDA ODA PDA CDA
w/ $L_{src}$ 85.2 86.7 85.3 80.0
Ours (w/o $L_{src}$) 86.8 86.9 84.9 77.0
Trade-off for Training Using the Source Samples. We observe that incorporating the source loss term $L_{src}$ introduces a tradeoff. Tab. 12 demonstrates that integrating source training significantly boosts performance in partial domain adaptation (PDA) and closed-set domain adaptation (CDA) compared to self-training. However, this approach leads to higher computational demands and slightly reduced performance in open-set domain adaptation (OPDA) and open domain adaptation (ODA). Therefore, the method can be selectively employed based on the objective (Tab. 12).
6 Limitations
A key limitation of our approach is that it does not explicitly address source-private class detection, which may result in suboptimal performance within the universal label space. However, we note that explicit source-private class detection often leads to mis-detection in practice, and focusing on target-private class detection can achieve most of the performance gains.
7 Conclusion
In this paper, we introduce Training-free Label Space Alignment (TLSA), a framework that leverages CLIP to address the Universal Domain Adaptation (UniDA) problem through label space alignment. TLSA utilizes a generative vision-language model to discover labels for all target domain samples and filters out those classified as source classes. On a mini-batch level, we eliminate lexical synonyms of source labels through synonym label alignment and remove semantically ambiguous labels via semantic label alignment. At the dataset level, frequency-based noisy candidate filtering removes unreliable labels based on occurrence patterns. This approach is highly efficient in both time and memory, completing the entire process in a single epoch without requiring further tuning. After alignment, we extend the source class set with the newly predicted target-private labels to construct a universal classifier, which significantly outperforms existing methods. Additionally, through self-training with balanced top-k confident sample selection, we surpass state-of-the-art UniDA baselines.
Data Availability
All the data utilized in this paper is publicly available at (Saenko et al., 2010; Venkateswara et al., 2017; Peng et al., 2017, 2019).
Declaration of generative AI in scientific writing
During the preparation of this work, the author used ChatGPT (Achiam et al., 2023) in order to improve readability. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
Appendix A Pseudo Code
In this section, we provide a pseudo-code of our semantic label alignment approach in Alg. 1.
Algorithm 1 Semantic Label Alignment
1:Target samples $\mathbf{x}^{t}$, filtered discovered labels $\tilde{R}$, source labels $C_{s}$
2:Frequency bank $F$
3:Compute target embeddings: $\mathbf{z}_{v}\leftarrow f_{v}(\mathbf{x}^{t})$
[…]
5:Compute similarity matrix: $S\leftarrow sim(\mathbf{z}_{v},\mathbf{z}_{t})$
6:Extract top-$k$ scores and indices: $(S_{k},I_{k})\leftarrow\operatorname{topk}(S)$
7:Find largest gap index: $\mathbf{J}\leftarrow\arg\max_{j}(s_{j+1}-s_{j})$
8:Define thresholds:
9:    $\tau_{\text{gap}}\leftarrow s_{\mathbf{J}}$
10:    $\tau_{\text{avg}}\leftarrow\mathbb{E}[S_{k}]$
11:    $\tau_{\text{set}}\leftarrow\min\{\tau_{\text{gap}},\tau_{\text{avg}}\}$
12:for $i=1$ to NN do
13:  Prediction set: $\mathcal{C}^{i}\leftarrow\{\mathbf{t}_{i,j}\mid j\in I_{k}^{i},\;s_{i,j}>\tau_{\text{set}}^{i}\}$
14:  if $\mathcal{C}^{i}\cap C_{s}=\emptyset$ then
15:   $F[r_{i}]\leftarrow F[r_{i}]+1$
16:  else
17:   $F[p_{i}]\leftarrow F[p_{i}]+1$
18:  end if
19:end for
Figure 10: Illustrations of two naive filtering methods: (a) shows the misalignment in margin-based alignment. The example demonstrates a case where $\tau$ is not suitable for this particular dataset. (b) illustrates the misalignment in top-1 prediction-based alignment. This example shows how misalignment occurs when the source label is slightly less similar to the discovered label
Appendix B Examples of Label Space Alignment
Fig. 10 illustrates the detailed process of two naive filtering methods, margin-based alignment and top-1 prediction-based alignment. In margin-based alignment, deriving the margin threshold $\tau$ is a non-trivial task because the class similarity between the source class and the target private class is not consistent across datasets (e.g., the case of “clipboard” in source labels and “notebook” in target private labels). In this scenario, “notebook” is consistently hindered by ”clipboard” due to the lack of a significantly high margin of similarity. Therefore, $\tau$ must be adaptively adjusted. Fig. 10a demonstrates this misalignment case.
Additionally, in top-1 prediction-based alignment, in the case of “backpack” in source labels and “bag” in target private labels, since “backpack” and “bag” are not synonyms according to the WordNet hierarchy, “bag” is often chosen as the target private candidate label on a per-instance basis. Fig. 10b demonstrates this misalignment case.
Figure 11: Label space alignment with examples. This figure demonstrates the alignment process by showcasing the discovered labels, corresponding candidate prediction sets, and the final filtered results for each item. The output is stored in the frequency bank, with counts incremented by one. Discovered labels are colored in violet, and source labels are colored in orange.
(a) An Example of Source Class Samples
GT backpack projector monitor
Image Refer to caption Refer to caption Refer to caption
Discovered label bag camera tv
Prediction set 1. backpack 1. projector 1. desktop
$\mathcal{C}^{i}$ 2. bag 2. camera 2. monitor
Output backpack projector monitor
(b) An Example of Target-Private Class Samples
GT printer stapler trash can
Image Refer to caption Refer to caption Refer to caption
Discovered label printer stapler recycling
Prediction set 1. fax machine 1. stapler 1. recycling
$\mathcal{C}^{i}$ 2. modem 2. trash can
3. recycle bin
Output printer stapler recycling
Appendix C Details of Semantic Label Alignment
Fig. 11 presents random examples from the Office31 dataset under the DSLR OPDA setting. It shows the exact discovered labels $r_{i}$ for each image $i$, the corresponding prediction set $\mathcal{C}^{i}$, and the specific elements stored in the frequency bank for each case. In the prediction set, the numbers indicate the ranking based on image-text similarity.
Appendix D Analysis of Labels to be Added to the Frequency Bank
In Sec. 3.4, we noted that storing BLIP’s discovered label to the frequency bank is more beneficial for performance compared to storing CLIP’s top-1 prediction to the frequency bank. As shown in Tab. 13, storing CLIP’s top-1 predictions often leads to an overestimation of incorrect target private labels (e.g., overly specific or entirely incorrect classes), which in turn causes the correct target private labels (e.g., ‘train’, ‘truck’, ‘skateboard’) to be underestimated. This imbalance ultimately results in degraded model performance. In Tab. 13, the correct target private label is underestimated when using CLIP’s top-1 prediction compared to using BLIP’s discovered label.
Table 13: Comparison of predicted target private labels between label selection strategies in label space alignment in VisDA open-partial setting. When BLIP’s discovered label is incorporated into the target private candidate labels, the H-score (%) is estimated at 86.3. In contrast, using CLIP’s top-1 prediction for target private candidate labels results in an H-score (%) of 80.1.
Ours (BLIP’s prediction) CLIP’s Top-1 Prediction
Label Frequency Label Frequency
Correct Target-Private Label
train 3328 train 3099
truck 2215 truck 1567
skateboard 2035 skateboard 1857
Incorrect Target-Private Label
fire truck 221 fire truck 293
van 533 van 574
dog 152 dog 96
dump truck 97 dump truck 205
skateboarder 108 skateboarder 350
food truck 136 food truck 214
pizza 113 pizza 93
umbrella 121 umbrella 87
taxi 172 taxi 388
parking meter 136 parking meter 252
jeep 113 jeep 105
fire hydrant 81 fork 94
carriage 96
locomotive 98
suv 110
police 83
freightliner 83
caboose 107
trolley 83
Appendix E Details of Prompts to Generative VLM
This section provides details of the question prompts used for label discovery. To ensure diversity while maintaining the same meaning, we employed five different styles of prompts. Tab. 14 demonstrates the detailed examples of question prompts for generative VLM.
Table 14: Examples of prompts used for generative VLM.
“What is the exact name of the object in the painting? Please tell me.”
“What can you identify in the illustration? Please tell me in a word.”
“What is the specific name of the object captured in the photograph? Please tell me.”
“What appears in the picture with its accurate name in the drawing? Please tell me.”
“What is portrayed in the photo with its exact name? Please tell me.”
Appendix F Time Complexity of Our Method
Fig. 12 illustrates the time complexity of our approach. Although two VLMs are employed, the framework is entirely training-free, introducing minimal overhead compared to prior works. Furthermore, it achieves faster adaptation than clustering-based methods (Chang et al., 2022).
A potential limitation lies in scalability, since generative VLM inference is required for each target-domain sample, leading to a runtime that grows linearly with dataset size. In practice, however, the cost remains manageable: even on DomainNet, the largest dataset in DomainBed, inference completed within 1.5 hours on a single GPU, demonstrating practical feasibility.
Figure 12: Average time spent on adaptation.
Appendix G Details of Main Results
This section details the experimental results (H-score, H3-score) for domain-to-domain. Tab. 16 through Tab. 26 offer a comprehensive comparison between baseline methods and the proposed approach using the CLIP ViT-B/16 backbone.
Table 15: Office31 ViT-B/16 OPDA
Methods A2D A2W D2A D2W W2A W2D Avg H-score SO 84.4 79.0 88.1 93.3 80.2 89.8 85.8 DANCE 91.6 84.7 93.7 95.8 87.8 96.4 91.7 OVANet 86.8 76.1 90.1 94.9 91.0 95.2 89.0 UniOT 78.5 73.4 82.5 88.4 83.0 88.1 82.3 CLIP-Distill 81.6 81.1 87.7 81.0 87.7 81.8 83.5 MemSPM+DCC 88.0 84.6 88.7 87.6 87.9 94.3 88.5 TLSA 92.7 86.8 89.7 86.8 89.7 92.7 89.9 TLSA+ST 95.4 88.7 91.5 88.7 91.5 95.4 91.9 H3-score SO 85.6 81.7 78.5 91.4 74.1 89.2 83.4 DANCE 90.4 85.7 81.4 93.0 78.3 93.4 87.0 OVANet 87.2 79.6 79.5 92.4 80.0 92.6 85.2 UniOT 80.5 76.9 75.2 88.5 75.4 87.7 80.7 CLIP-Distill 83.7 83.2 78.2 83.2 78.2 83.8 81.7 TLSA 91.1 87.1 79.4 87.1 79.4 91.1 85.9 TLSA+ST 88.7 85.2 81.9 85.2 81.9 88.7 85.3
Table 16: Office31 ViT-B/16 ODA
Methods A2D A2W D2A D2W W2A W2D Avg H-score SO 87.2 87.9 85.1 90.7 84.1 86.0 86.8 DANCE 92.4 89.3 93.4 96.7 91.7 95.6 93.2 OVANet 86.7 80.4 91.7 97.4 84.4 95.1 89.3 UniOT 91.4 91.0 92.4 96.8 94.2 97.7 93.9 CLIP-Distill 80.1 79.7 88.0 81.2 90.2 86.5 84.3 TLSA 92.2 91.7 94.1 91.7 94.1 92.2 92.7 TLSA+Self training 94.3 95.4 95.9 95.4 95.9 94.3 95.2 H3-score SO 87.8 88.8 80.5 90.6 79.9 87.0 85.8 DANCE 91.3 89.7 85.3 94.6 84.3 93.4 89.8 OVANet 87.5 83.5 84.4 95.0 80.1 93.1 87.3 UniOT 90.7 89.7 83.8 94.2 85.1 95.5 89.8 CLIP-Distill 82.9 83.0 82.2 84.1 83.4 87.4 83.8 TLSA 91.2 91.1 85.7 91.1 85.7 91.2 89.3 TLSA+ST 90.9 91.6 87.1 91.6 87.1 90.9 89.9
Table 17: Office31 ViT-B/16 PDA
Methods A2D A2W D2A D2W W2A W2D Avg H-score SO 77.0 67.6 87.2 96.9 78.5 95.4 83.7 DANCE 51.7 49.3 86.2 89.3 67.2 99.0 73.8 OVANet 74.6 64.4 87.7 97.2 82.7 97.6 84.0 UniOT 28.9 26.4 27.8 38.1 29.5 38.3 31.5 CLIP-Distill 82.6 82.7 91.5 82.3 92.1 82.6 85.6 TLSA 92.7 86.8 89.7 86.8 89.7 92.7 89.9 TLSA+ST 96.2 98.4 96.1 98.4 96.1 96.2 96.9
Table 18: Office31 ViT-B/16 CDA
Methods A2D A2W D2A D2W W2A W2D Avg H-score SO 63.4 57.5 58.1 93.3 56.1 96.2 70.8 DANCE 58.5 50.7 53.7 93.3 52.3 97.6 67.7 OVANet 56.5 51.8 51.3 89.1 51.2 95.7 65.9 UniOT 51.9 49.8 48.4 65.3 46.2 63.7 54.2 CLIP-Distill 75.5 72.0 73.6 71.4 74.5 75.4 73.7 TLSA 82.7 80.5 79.4 80.5 79.4 82.7 79.0 TLSA+ST 84.6 84.0 80.1 84.0 80.1 84.6 82.9
Table 19: Office-Home ViT-B/16 OPDA
Methods A2C A2P A2R C2A C2P C2R P2A P2C P2R R2A R2C R2P Avg H-score SO 69.8 80.3 83.8 80.5 81.5 85.3 77.5 71.2 84.7 80.6 76.0 77.6 79.0 DANCE 78.0 86.4 87.1 81.1 85.2 85.6 82.4 79.9 90.5 83.6 79.6 84.1 83.6 OVANet 72.7 88.0 86.4 80.6 84.0 85.3 75.2 63.6 89.1 81.8 74.4 85.0 80.5 UniOT 85.5 89.8 91.1 79.9 87.3 89.3 80.8 83.5 88.9 83.5 84.3 91.6 86.3 CLIP-Distill 81.4 79.6 86.1 87.9 80.3 86.7 87.0 81.3 91.0 87.0 81.9 86.8 84.7 MemSPM+DCC 78.1 90.3 90.7 81.9 90.5 88.3 79.2 77.4 87.8 78.8 76.2 91.6 84.2 TLSA 87.4 95.2 95.8 86.8 95.2 95.8 86.8 87.4 95.8 86.8 87.4 95.2 91.3 TLSA+ST 90.2 96.3 96.7 90.6 96.3 96.7 90.6 90.2 96.7 90.6 90.2 96.3 93.4 H3-score SO 67.5 82.4 82.7 73.7 83.2 83.7 71.9 68.4 83.3 73.7 71.3 80.5 76.9 DANCE 72.4 86.6 84.8 74.0 85.8 83.8 74.7 73.5 86.9 75.4 73.3 85.0 79.7 OVANet 69.3 87.7 84.4 73.7 85.0 83.7 70.6 63.5 86.1 74.4 70.3 85.7 77.9 UniOT 78.9 89.4 88.1 75.9 87.9 87.1 76.2 76.9 86.6 77.8 77.9 90.5 82.8 CLIP-Distill 74.4 81.9 84.2 77.6 82.4 84.6 77.1 74.3 87.3 77.1 74.6 86.9 80.2 TLSA 79.8 92.2 89.5 76.8 92.7 89.5 76.8 79.8 89.5 76.8 79.8 92.2 84.6 TLSA+ST 85.5 93.9 94.1 87.5 93.9 94.1 87.5 85.5 94.1 87.5 85.5 93.9 90.2
Table 20: Office-Home ViT-B/16 ODA
Methods A2C A2P A2R C2A C2P C2R P2A P2C P2R R2A R2C R2P Avg H-score SO 69.6 77.3 80.5 76.8 81.0 82.2 70.2 69.7 81.4 78.7 73.2 77.1 76.5 DANCE 75.2 81.5 82.9 76.2 77.1 81.6 70.2 71.5 85.6 80.0 74.8 82.6 78.3 OVANet 71.6 81.7 82.7 76.5 83.3 81.6 64.0 56.9 84.1 78.6 70.5 84.4 76.3 UniOT 81.0 85.9 85.0 76.8 85.5 83.9 72.1 77.8 83.7 80.4 79.3 89.7 81.7 CLIP-Distill 77.3 79.1 84.6 83.4 79.7 85.1 79.6 75.9 88.6 79.7 76.9 85.9 81.3 MemSPM+DCC 69.7 83.2 85.2 72.0 79.2 81.2 72.3 66.7 85.2 72.7 66.0 84.5 76.5 TLSA 85.5 92.3 93.9 85.1 92.3 93.9 85.1 85.5 93.9 85.1 85.5 92.3 89.2 TLSA+ST 89.2 94.4 94.2 88.0 94.4 94.2 88.0 89.2 94.2 88.0 89.2 94.4 91.4 H3-score SO 67.4 80.3 80.5 71.5 82.9 81.7 67.6 67.5 81.2 72.6 69.6 80.2 75.2 DANCE 70.8 83.3 82.1 71.2 80.1 81.3 67.6 68.6 83.9 73.4 70.5 84.1 76.4 OVANet 68.7 83.4 82.0 71.4 84.5 81.3 63.7 58.9 82.9 72.6 68.0 85.2 75.2 UniOT 75.7 86.6 83.9 73.4 86.8 83.3 70.2 73.7 83.1 76.0 74.6 89.1 79.7 CLIP-Distill 72.0 81.6 83.3 75.2 82.0 83.6 73.1 71.2 85.8 73.2 71.8 86.3 78.3 TLSA 78.7 90.4 88.3 75.9 90.4 88.3 75.9 78.7 88.3 75.9 78.7 90.4 83.3 TLSA+ST 85.0 92.8 92.3 85.7 92.8 92.3 85.7 85.0 92.3 85.7 85.0 92.8 89.0
Table 21: Office-Home ViT-B/16 PDA
Methods A2C A2P A2R C2A C2P C2R P2A P2C P2R R2A R2C R2P Avg H-score SO 37.8 58.5 63.7 38.6 56.4 59.4 41.0 32.9 68.5 50.2 43.4 71.1 51.8 DANCE 34.3 52.6 66.0 30.2 41.5 59.1 30.7 19.9 68.2 47.9 39.5 72.4 46.8 OVANet 51.5 62.7 73.1 52.3 59.4 65.7 44.8 35.4 73.3 61.7 49.1 76.8 58.8 UniOT 31.8 34.3 31.1 26.0 32.2 29.4 26.9 30.0 33.1 36.9 31.5 38.4 31.8 CLIP-Distill 64.9 86.7 87.6 71.5 85.6 86.5 63.3 57.2 82.9 65.7 59.0 82.5 74.4 MemSPM+DCC 64.7 81.1 84.5 74.8 74.7 77.5 58.7 60.3 84.2 70.3 77.2 85.8 74.5 TLSA 74.5 87.8 90.4 75.3 87.8 90.4 75.3 74.5 90.4 75.3 74.5 87.8 82.0 TLSA+ST 79.2 86.3 93.5 82.2 86.3 93.5 82.2 79.2 93.5 82.2 79.2 86.3 85.3
Table 22: Office-Home ViT-B/16 CDA
Methods A2C A2P A2R C2A C2P C2R P2A P2C P2R R2A R2C R2P Avg H-score SO 31.3 57.0 59.8 36.8 60.9 56.5 35.6 28.3 64.9 50.2 36.3 77.5 49.6 DANCE 31.1 57.1 60.6 36.7 58.3 56.7 33.1 28.5 64.6 50.5 36.2 78.8 49.4 OVANet 41.5 60.5 70.1 49.8 61.6 64.2 42.3 30.7 70.4 60.2 40.8 81.5 56.1 UniOT 41.4 46.9 50.9 41.7 50.1 50.6 36.8 38.0 50.0 46.2 42.9 61.5 46.4 CLIP-Distill 53.1 85.9 83.2 65.1 84.9 82.6 59.2 45.7 78.5 60.4 47.8 82.2 69.0 TLSA 66.4 87.2 88.2 74.2 87.2 88.2 74.2 66.4 88.2 74.2 66.4 87.2 79.0 TLSA+ST 71.3 88.6 89.9 77.3 88.6 89.9 77.3 71.3 89.9 77.3 71.3 88.6 81.8
Table 23: DomainNet ViT-B/16 OPDA
Methods P2R P2S R2P R2S S2P S2R Avg H-score SO 64.2 50.2 53.1 52.1 48.3 67.1 55.8 DANCE 64.2 50.1 50.8 51.2 47.4 67.2 55.1 OVANet 68.2 58.6 60.5 59.1 56.9 68.6 62.0 UniOT 71.4 62.8 62.5 64.6 59.5 72.7 65.6 CLIP-Distill 74.1 61.6 58.5 60.2 59.7 74.8 64.8 MemSPM+DCC 62.4 52.8 58.5 53.3 50.4 62.6 56.7 TLSA 83.6 71.2 71.3 71.2 71.3 83.6 75.4 TLSA+ST 82.3 69.4 71.2 69.4 71.2 82.3 74.3 H3-score SO 69.3 52.5 57.2 53.8 53.4 71.5 59.6 DANCE 69.3 52.4 55.4 53.2 52.6 71.5 59.1 OVANet 72.3 58.3 62.7 58.6 60.1 72.6 64.1 UniOT 73.9 61.6 64.0 63.3 61.6 75.1 66.6 CLIP-Distill 76.6 60.2 61.2 59.4 62.1 77.1 66.1 TLSA 83.1 69.8 66.1 69.8 66.1 83.1 73.1 TLSA+ST 84.0 69.2 72.9 69.2 72.9 84.0 75.4
Table 24: DomainNet ViT-B/16 ODA
Methods P2R P2S R2P R2S S2P S2R Avg H-score SO 68.5 55.0 57.0 56.6 51.3 70.0 59.7 DANCE 68.6 54.7 55.5 56.3 51.1 70.2 59.4 OVANet 70.4 61.4 62.5 62.1 59.4 70.6 64.4 UniOT 76.1 66.9 66.8 68.2 63.3 75.3 69.4 CLIP-Distill 76.2 63.6 59.7 62.4 61.6 76.5 66.7 TLSA 85.6 72.9 72.1 72.9 72.1 85.6 76.9 TLSA+ST 83.1 70.4 71.2 70.4 71.2 83.1 74.9 H3-score SO 72.2 56.6 59.6 57.8 55.3 73.3 62.5 DANCE 72.3 56.5 58.5 57.6 55.2 73.4 62.2 OVANet 73.7 61.0 63.4 61.5 61.3 73.8 65.8 UniOT 76.7 65.0 66.3 66.3 63.6 76.3 69.0 CLIP-Distill 77.7 62.5 61.5 61.7 62.8 77.9 67.4 TLSA 83.7 68.1 69.5 68.1 69.5 83.7 73.8 TLSA+ST 83.9 70.3 72.3 70.3 72.3 83.9 75.5
Table 25: DomainNet ViT-B/16 PDA
Methods P2R P2S R2P R2S S2P S2R Avg H-score SO 44.4 26.5 30.0 28.6 24.0 44.0 32.9 DANCE 43.2 25.3 26.3 25.4 20.2 39.8 30.0 OVANet 55.8 36.9 41.1 39.1 38.6 58.4 45.0 UniOT 46.5 39.4 39.1 41.4 33.6 45.2 40.9 CLIP-Distill 85.0 75.0 49.1 47.9 46.4 75.0 63.1 TLSA 84.6 64.7 60.9 64.7 60.9 84.6 70.1 TLSA+ST 86.8 70.0 66.4 70.0 66.4 86.8 74.4
Table 26: DomainNet ViT-B/16 CDA
Methods P2R P2S R2P R2S S2P S2R Avg H-score SO 46.9 26.6 30.4 29.3 26.7 48.8 34.8 DANCE 46.7 26.5 29.7 28.3 26.1 47.1 34.1 OVANet 57.6 36.9 40.9 39.7 41.2 60.8 46.2 UniOT 60.0 47.1 46.8 49.7 44.6 60.9 51.5 CLIP-Distill 71.3 42.9 39.5 42.3 40.2 71.5 51.3 TLSA 83.7 62.6 60.3 62.6 60.3 83.7 68.9 TLSA+ST 81.0 61.6 61.1 61.6 61.1 81.0 67.9